OpenStudy (anonymous):

Find the slope of a line passing through (3, -9) and (6, -1). A. three-eighths B. negative three-eighths C. eight-thirds D. negative eight-thirds

OpenStudy (anonymous):

c

OpenStudy (anonymous):

m=(Y2-Y1)/(X2-X1)

OpenStudy (anonymous):

-1 + 9 / 6-3 = 8 / 3 c

OpenStudy (anonymous):

Step 1. m = (-1-(-9))/(6-3) Step 2. m= 8/3